# DiffMap: Enhancing Map Segmentation with Map Prior Using Diffusion Model

### v1

@@ Abstract
Constructing high-definition (HD) maps is a crucial requirement for enabling autonomous driving. In recent years, several map segmentation algorithms have been developed to address this need, leveraging advancements in Bird’s-Eye View (BEV) perception. However, existing models still encounter challenges in producing realistic and consistent semantic map layouts. One prominent issue is the limited utilization of structured priors inherent in map segmentation masks. In light of this, we propose DiffMap, a novel approach specifically designed to model the structured priors of map segmentation masks using latent diffusion model. By incorporating this technique, the performance of existing semantic segmentation methods can be significantly enhanced and certain structural errors present in the segmentation outputs can be effectively rectified. Notably, the proposed module can be seamlessly integrated into any map segmentation model, thereby augmenting its capability to accurately delineate semantic information. Furthermore, through extensive visualization analysis, our model demonstrates superior proficiency in generating results that more accurately reflect real-world map layouts, further validating its efficacy in improving the quality of the generated maps.
@@ I Introduction
HD maps are essential for facilitating accurate environmental understanding and precise navigation in autonomous vehicles. However, the manual construction of these maps is a laborious and complex process. To address this challenge, the integration of map construction into the BEV (Bird’s-Eye View) perception task has gained considerable attention. Current research regards the construction of rasterized HD maps as a segmentation task in the BEV space, primarily involving architectures similar to FCNs[1] (Fully Convolutional Networks) with segmentation heads after obtaining the BEV features. For instance, in short-range perception, HDMapNet[2] encodes sensor features through LSS[3] and then employs multi-branch FCNs for semantic segmentation, instance detection, and direction prediction to construct the map. For long-range map construction, SuperFusion[4] follows the similar pipeline with camera and LiDAR fusion mechanisms. Despite ongoing improvements in map segmentation, the pixel-based classification approach of segmentation exhibits inherent limitations, including the potential neglect of specific category attributes, which can lead to issues such as distorted and interrupted dividers, blurred pedestrian crossings, and other types of artifacts and noise, as shown in Fig.1 (a). These problems not only affect the structural accuracy of the map but could also directly impact downstream path planning modules of autonomous driving systems.
@@
Therefore, we propose to use a modified LDM as an enhancement module to model map structured prior for existing segmentation models. We design a network named DiffMap, which not only learns the map prior through a process of adding and removing noise, but also integrates BEV features from classic BEV feature extraction networks as control signals to ensure that the outputs match the current frame observations. Our experimental results demonstrate that this module can effectively generate a more smooth and reasonable map segmentation while greatly reduce the artifacts, therefore improving the overall map construction performance. This is the first study to incorporate a LDM model into map prior modeling and use as a plug-and-play module to enhance any map segmentation task. We plan to share all our code and data to facilitate further research and application in the community.
@@ II-A Semantic map construction
In traditional HD maps construction, semantic maps are typically manually or semi-automatically annotated based on the LiDAR point clouds. This involves using SLAM-based algorithms[9, 10] to build globally consistent maps, with semantic annotations manually added to the maps. However, this approach is time-consuming and labor-intensive, and it presents challenges in terms of updating the maps, thereby limiting scalability and real-time performance. HDMapNet[2] pioneered a method for dynamically constructing local semantic maps using onboard sensors. It encodes LiDAR point clouds and surround-view image features into BEV space, and decodes with three different heads. Finally produces a vectorized local semantic map with post-processing. . SuperFusion[4], in contrast, focuses on constructing long-range high-precision semantic maps, leveraging LiDAR depth for enhancing image depth estimation and employing image features to guide remote LiDAR feature prediction. And then employs a map detection head akin to HDMapNet to get semantic maps. MachMap[11] divides the task into line detection of polylines and instance segmentation of polygons and uses postprocessing to refine the mask to obtain the final result. The subsequent work[12, 13, 14, 15] explores end-to-end online mapping and gets the direct vectorized HD map. Dynamic construction of semantic maps without manual annotation effectively reduces the construction cost.
@@ II-B Diffusion model for segmentation and detection
Denoising Diffusion Probabilistic Models (DDPMs) [16] is a class of generative models based on Markov chains, which has demonstrated outstanding performance in fields such as image generation. Nowadays, DDPMs has gradually expanded its application to various tasks including segmentation and detection. SegDiff[17] applies the diffusion model to address image segmentation tasks, where the UNet encoder used in the diffusion model is further decoupled into three modules: E, F, and G. Modules G and F encode the input image $\mathbf{I}$ and segmentation map $\mathbf{X}_{t}$, respectively, which are then merged by addition in E to iteratively refine the segmentation map. DDPMS[6] produces an initial prediction prior with base segmentation models, and refines the prior with diffusion model. DiffusionDet[18] extends the diffusion model to the object detection framework, modeling object detection as a denoising diffusion process from noisy boxes to object boxes. Diffusion is also utilized in the field of autonomous driving, such as in the case of MagicDrive[19], where it is applied in the synthesis of street views guided by geometric constraints and Motiondiffuser[20] extends the diffusion model to the multi-agent motion prediction problems.
@@ III-A Preliminaries
DDPMs are a type of generative models that leverage diffusion processes to approximate the underlying data distribution $p(\mathbf{x})$. They function by systematically removing noise from a variable that is normally distributed, echoing the reverse operation of a Markov Chain of predetermined length $T$. Within DDPMs, the transformation from a clean initial data distribution $\mathbf{x}_{0}$ to a noisy dataset $\mathbf{x}_{T}$ is accomplished through a series of conditional probabilities $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$, with $q$ denoting the noise introduction process. At each step $t$, noise is added according to the following specification:
@@
Through this formulation, the LDMs’ denoising core, represented by $\epsilon_{\theta}$, is instantiated as a time-conditioned UNet. With the noise induction process being deterministic, $\mathbf{z}_{t}$ is efficiently obtained from $\epsilon$ during training, and the model is capable of generating images from $p(\mathbf{z})$ in a single forward pass through the network $D$. This approach not only enhances the image synthesis efficiency but also ensures the production of high-fidelity images with significantly lower computational costs during both training and inference stages.
@@
This configuration ensures that $\tau_{\theta}$ and $\epsilon_{\theta}$ are jointly optimized, showcasing the adaptability of the conditioning mechanism. Such a system can be enriched with specialized knowledge for various domains, for example, by incorporating transformer models for text-based prompts.
@@ III-B Architecture
We propose DiffMap as a decoder to incorporate the diffusion model into semantic map segmentation models. The overall framework of DiffMap is illustrated in Fig. 2. The model takes surrounding multi-view images and LiDAR point clouds as inputs, encodes them into BEV space and get the fused BEV features. DiffMap is then employed as the decoder to generate segmentation maps. Within the DiffMap module, we utilize the BEV features as conditions to guide the denoising process. In this section, we first introduce the base model for generating BEV features, followed by an explanation of DiffMap and its denoising module. Finally, we present the training and inference processes.
Figure 2: Architecture Overview: After extracting features from surrounding multi-view images and LiDAR point clouds separately using backbone networks, the features are transformed into Bird’s Eye View (BEV) space for fusion. During the training process, random noise is continuously added to the ground truth map. Then, in the denoising process, the fused BEV features are used as conditional control variables of Diffmap, ultimately generating the predicted segmentation map. Whereas in inference process, results are obtained in the continuous denoising from random noise.
@@ III-B2 DiffMap Module
Firstly, We add noise $\sigma$ through diffusion process and get the noisy latent map $\left\{\mathbf{z}_{t}\right\}_{t=0}^{T}$ at each timestep $t$, where $\mathbf{z}_{0}=\mathbf{z}$. Then during the denoising process, UNet serves as the backbone network for the noise prediction. Given that our primary objective is to utilize the diffusion model to refine map segmentation results, we hope to enhance the supervision of the segmentation results and expect the DiffMap model to directly provide semantic features for instance-related predictions during training. Thus, we further decouple the UNet network structure into two branches, one branch is to predict noise $\epsilon$ as traditional diffusion model and the other is to predict $\mathbf{z}$ in the latent space. The overall denoising module is illustrated in Fig. 3. After getting the latent map prediction $\mathbf{z}_{\theta}$, we decode it to the original pixel space as the semantic feature map. Then we can obtain the instance prediction from it following the method proposed by HDMapNet, which outputs three predictions with different heads: semantic segmentation, instance embedding, and lane direction. These predictions are subsequently used in the post-processing step to vectorize the map.
Figure 3: Denoising Module: In order to get a direct result of the segmented features, we decouple the UNet decoder into two branches, one branch is to predict noise $\epsilon$ as traditional diffusion model and the other is to predict the latent groundtruth $z_{0}$. During the denoising process, we resize BEV features into latent space size as a conditional control variable. We first concatenate it with noisy latent map, and then incorporate it into the two decoders of UNet network with cross attention mechanism.
@@
Secondly, we incorporate a cross attention mechanism into the each layer of UNet network, where $\mathcal{B}$ plays as key/value and $\mathbf{z}_{t}$ plays as query, as shown in Fig. 3. The formulation of the cross attention module is illustrated as:
@@ III-C1 Training
where $z_{e}(\mathbf{x})$ stands for the output of the encoder, $e$ stands for the embedding space, $z_{q}(\mathbf{x})$ stands for the the nearest embedding of $z_{e}(\mathbf{x})$ in the embedding space, and sg is the stop-gradient operator. Through VQVAE, the ground truth maps can be compressed into the latent space, improving generation efficiency and preventing the diffusion model from over-training on image pixels.
@@
In the equation, ${\mathbf{z}}_{\theta}$ and ${\epsilon}_{\theta}$ correspond to the image branch and noise branch, respectively. Subsequently, the model is trained to predict images and noise separately. Additionally, to achieve higher-level image generation accuracy, we also incorporate the losses of baseline model as auxiliary supervision, including the cross-entropy loss for the semantic segmentation, and the discriminative loss for the instance embedding mentioned in HDMapNet. The final objective function combines the loss of diffusion model $\mathcal{L}_{\text{diff}}$ with the auxiliary losses from the baseline model $\mathcal{L}_{\text{baseline}}$, resulting in the following mathematical expression:
@@ III-C2 Inference
As for the sampling in the denoising iterations, we adopt DDIM sampler[29] to accelerate sampling. Specifically, we can generate a sample $\mathbf{x}_{t-1}$ from a sample $\mathbf{x}_{t}$ via:
$\mathbf{x}_{t-1}=$ $\sqrt{\alpha_{t-1}}\underbrace{\left(\frac{\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}^{(t)}\left(\mathbf{x}_{t}\right)}{\sqrt{\alpha_{t}}}\right)}_{\text{predicted }\bm{x}_{0}}$ (10)
$+\underbrace{\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot\epsilon_{\theta}^{(t)}\left(\mathbf{x}_{t}\right)}_{\text{direction pointing to }\mathbf{x}_{t}}+\underbrace{\sigma_{t}\epsilon_{t}}_{\text{random noise}}$
where $\epsilon_{t}\sim\mathcal{N}(\mathbf{0},\bm{I})$ is standard Gaussian noise independent of $\mathbf{x}_{t}$. $\sigma_{t}$ is a hyper-parameter, and different choices of $\sigma_{t}$ values results in different generative processes. When $\sigma_{t}=0$ for all $t$, the forward process becomes deterministic given $\mathbf{x}_{t-1}$ and $\mathbf{x}_{0}$, except for $t=1$; in the generative process, the coefficient before the random noise $\epsilon_{t}$ becomes zero. The term $\alpha_{t}$ is signal rate, equalling $1-\beta_{t}$ mentioned in Equation 1.
@@ IV-A Implementation Details
(+5.5) (+1.1) (+2.2) (+6.3) (+3.3) (+1.9)
@@ IV-A1 Dataset
We validate our DiffMap on the nuScenes dataset[30]. As a popular benchmark in autonomous driving, the nuScenes dataset includes multi-view images and point cloud of 1,000 scenes, with 700 scenes allocated for training, 150 for validation, and 150 for testing. The nuScenes dataset also has annotated HD map semantic labels, which are highly suitable for our task as well. Besides training the main model, we pre-processed the nuScenes dataset into segmentation maps during the pre-training of the autoencoder.
@@ IV-A2 Architecture
We use ResNet-101[31] as the backbone for our camera branch and PointPillars[32] as our LiDAR branch backbone. The Segmentation head in baseline is a FCN network based ResNet-18. For the autoencoder, we adpot VQVAE, which is pre-trained on the nuScenes segmentation maps dataset, to extract map features and compress map into the basic latent space. UNet is used to construct the diffusion network.

### v2

@@ Abstract
Constructing high-definition (HD) maps is a crucial requirement for enabling autonomous driving. In recent years, several map segmentation algorithms have been developed to address this need, leveraging advancements in Bird’s-Eye View (BEV) perception. However, existing models still encounter challenges in producing realistic and consistent semantic map layouts. A prominent issue is the limited utilization of structured priors inherent in map segmentation masks. In light of this, we propose DiffMap, a novel approach specifically designed to model the structured priors of map segmentation masks using latent diffusion model. By incorporating this technique, the performance of existing semantic segmentation methods can be significantly enhanced and certain structural errors present in the segmentation outputs can be effectively rectified. Notably, the proposed module can be seamlessly integrated into any map segmentation model, thereby augmenting its capability to accurately delineate semantic information. Furthermore, through extensive visualization analysis, our model demonstrates superior proficiency in generating results that more accurately reflect real-world map layouts, further validating its efficacy in improving the quality of the generated maps.
Index Terms:
Autonomous Vehicle Navigation, Sensor Fusion, Deep Learning for Visual Perception
@@ I Introduction
HD maps are essential for facilitating accurate environmental understanding and precise navigation in autonomous vehicles. However, the traditional generation process of these maps is a laborious and complex process. To address this challenge, the integration of map construction into the BEV (Bird’s-Eye View) perception task has gained considerable attention. Current research regards the construction of rasterized HD maps as a segmentation task in the BEV space, primarily involving architectures similar to Fully Convolutional Networks (FCNs)[1] with segmentation heads after obtaining the BEV features. For instance, in short-range perception, HDMapNet[2] encodes sensor features through LSS[3] and then employs multi-branch FCNs for semantic segmentation, instance detection, and direction prediction to construct the map. For long-range map construction, SuperFusion[4] follows the similar pipeline with camera and LiDAR fusion mechanisms. Despite ongoing improvements in map segmentation, the pixel-based classification approach of segmentation exhibits inherent limitations, including the potential neglect of specific category attributes, which can lead to issues such as distorted and interrupted dividers, blurred pedestrian crossings, and other types of artifacts and noise, as shown in Fig.1 (a). These problems not only affect the structural accuracy of the map but could also directly impact downstream path planning modules of autonomous driving systems.
@@
Therefore, we propose a network named DiffMap for online map reconstruction task. The network uses a modified LDM as an enhancement module to model map structured prior for existing segmentation models. To migrate the diffusion technique into the online mapping task, there are two main challenges need to be resolved. First, the mapping results are highly determined by online sensor observations. As a result, we design a condition injection module that integrates the online BEV feature of sensor observations into the denoising process as control signals. Second, since downstream task usually requires instance-wise mapping result, to enable the diffusion network to reconstruct the map elements separately, an additional discriminative module is proposed to help the network learn the distinguishable features of different map elements. Our experimental results demonstrate that DiffMap can effectively generate a more smooth and reasonable map segmentation while greatly reducing the artifacts, therefore improving the overall map construction performance. This is the first study to incorporate a LDM model into map prior modeling and be used as a plug-and-play module to enhance any map segmentation task. We plan to share all our code and data to facilitate further research and application in the community.
@@ II-A Semantic map construction
In traditional HD maps construction, semantic maps are typically manually or semi-automatically annotated based on the LiDAR point clouds. This involves using SLAM-based algorithms[9, 10] to build globally consistent maps, with semantic annotations manually added to the maps. However, this approach is time-consuming and labor-intensive, and it presents challenges in terms of updating the maps, thereby limiting scalability and real-time performance. HDMapNet[2] pioneered a method for dynamically constructing local semantic maps using onboard sensors. It encodes LiDAR point clouds and surround-view image features into BEV space, and decodes with three different heads. Finally produce a vectorized local semantic map with post-processing. SuperFusion[4], in contrast, focuses on constructing long-range high-precision semantic maps, leveraging LiDAR depth for enhancing image depth estimation and employing image features to guide remote LiDAR feature prediction. And then employs a map detection head akin to HDMapNet to get semantic maps. MachMap[11] divides the task into line detection of polylines and instance segmentation of polygons and uses post-processing to refine the mask to obtain the final result. The subsequent work[12, 13, 14, 15] explores end-to-end online mapping and gets the direct vectorized HD map. Dynamic construction of semantic maps without manual annotation effectively reduces the construction cost.
@@ II-B Diffusion model for segmentation and detection
Denoising Diffusion Probabilistic Models (DDPMs) [16] are a class of generative models based on Markov chains, which have demonstrated outstanding performance in fields such as image generation. Nowadays, DDPMs have gradually expanded their application to various tasks including segmentation and detection. SegDiff[17] applies the diffusion model to address image segmentation tasks, where the UNet encoder used in the diffusion model is further decoupled into three modules: E, F, and G. Modules G and F encode the input image $\mathbf{I}$ and segmentation map $\mathbf{X}_{t}$, respectively, which are then merged by addition in E to iteratively refine the segmentation map. DDPMS[6] produce an initial prediction prior with base segmentation models, and refines the prior with diffusion model. DiffusionDet[18] extends the diffusion model to the object detection framework, modeling object detection as a denoising diffusion process from noisy boxes to object boxes. Diffusion is also utilized in the field of autonomous driving, such as in the case of MagicDrive[19], where it is applied in the synthesis of street views guided by geometric constraints and Motiondiffuser[20] extends the diffusion model to the multi-agent motion prediction problems.
@@ III-A Preliminaries
DDPMs are a type of generative models that leverage diffusion processes to approximate the underlying data distribution $p(\mathbf{x})$. They function by systematically removing noise from a variable that is normally distributed, echoing the reverse operation of a Markov Chain of predetermined length $T$. Within DDPMs, the transformation from a clean initial data point $\mathbf{x}_{0}$ to a noisy one $\mathbf{x}_{T}$ is accomplished through a series of conditional probabilities $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$, with $q$ denoting the noise introduction process. At each step $t$, noise is added according to the following specification:
@@
where $\mathbf{z}_{t}=D(\mathbf{x}_{t})$ is the corresponding data point of $\mathbf{x}_{t}$ in the latent space. The transformation $D$ is modeled as a feedforward neural network. The denoising process of LDMs $\epsilon_{\theta}$ is instantiated as a time-conditioned UNet applied on $\mathbf{z}_{t}$. This approach not only enhances the image synthesis efficiency but also ensures the production of high-fidelity images with significantly lower computational costs during both training and inference stages.
@@
where $\tau_{\theta}$ is a domain specific encoder that projects condition $\mathbf{y}$ to an intermediate representation. This configuration ensures that $\tau_{\theta}$ and $\epsilon_{\theta}$ are jointly optimized, showcasing the adaptability of the conditioning mechanism. Such a system can be enriched with specialized knowledge for various domains, for example, by incorporating transformer models for text-based prompts.
@@ III-B Architecture
We propose DiffMap as a decoder to incorporate the diffusion model into semantic map segmentation models. The overall framework of DiffMap is illustrated in Fig. 2. The model takes surrounding multi-view images and LiDAR point clouds as inputs, encodes them into BEV space and gets the fused BEV features. DiffMap is then employed as the decoder to generate segmentation maps. Within the DiffMap module, we utilize the BEV features as conditions to guide the denoising process. In this section, we first introduce the base model for generating BEV features, followed by an explanation of DiffMap and its denoising module. Finally, we present the training and inference processes.
Figure 2: Architecture Overview: After extracting features from surrounding multi-view images and LiDAR point clouds separately using backbone networks, the features are transformed into Bird’s Eye View (BEV) space for fusion. During the training process, random noise is continuously added to the ground truth map. Then, in the denoising process, the fused BEV features are used as conditional control variables of Diffmap, ultimately generating the predicted segmentation map. Whereas in the inference process, results are obtained in the continuous denoising from random noise.
@@ III-B2 DiffMap Module
Firstly, We add noise $\sigma$ through diffusion process and get the noisy latent map $\left\{\mathbf{z}_{t}\right\}_{t=0}^{T}$ at each timestep $t$, where $\mathbf{z}_{0}=\mathbf{z}$. Then during the denoising process, UNet serves as the backbone network for the noise prediction. Given that our primary objective is to utilize the diffusion model to refine map segmentation results, we hope to enhance the supervision of the segmentation results and expect the DiffMap model to directly provide semantic features for instance predictions during training.
Thus, we further decouple the UNet network structure into two branches like [29], one branch is to predict noise $\epsilon$ as traditional diffusion model and the other is to predict $\mathbf{z}$ in the latent space. The overall denoising module is illustrated in Fig. 3.
@@
Secondly, we incorporate a cross attention mechanism into each layer of UNet, where $\mathcal{B}$ plays as key/value and $\mathbf{z}_{t}$ plays as query, as shown in Fig. 3. The formulation of the cross attention module is illustrated as:
@@
III-B3 Instance Segmentation
After getting the latent map prediction $\mathbf{z}_{\theta}$, we decode it to the original pixel space as the semantic feature map. Then we can obtain the instance prediction from it following the method proposed by HDMapNet, which outputs three predictions with different heads composed by MLPs: semantic segmentation, instance embedding, and lane direction. These predictions are subsequently used in the post-processing step to vectorize the map.
Figure 3: Denoising Module: In order to get a direct result of the segmented features, we decouple the UNet decoder into two branches, one branch is to predict noise $\epsilon$ as traditional diffusion model and the other is to predict the $\mathbf{z}$ in latent space. During the denoising process, we resize BEV features into latent space size as a conditional control variable. We first concatenate it with noisy latent map, and then incorporate it into the two decoders of UNet with cross attention mechanism.
@@ III-C1 Training
where $z_{e}(\mathbf{x})$ stands for the output of the encoder, $e$ stands for the embedding space, $z_{q}(\mathbf{x})$ stands for the the nearest embedding of $z_{e}(\mathbf{x})$ in the embedding space, and $sg(\cdot)$ is the stop-gradient operator. Through VQVAE, the ground truth maps can be compressed into the latent space, improving generation efficiency and preventing the diffusion model from over-training on image pixels.
@@
In the equation, ${\mathbf{z}}_{\theta}$ and ${\epsilon}_{\theta}$ correspond to the image branch and noise branch, respectively. Subsequently, the model is trained to predict images and noise separately. Additionally, to achieve higher-level image generation accuracy, we also incorporate the losses from baseline model as auxiliary supervision, including the cross-entropy loss for the semantic segmentation, and the discriminative loss for the instance embedding mentioned in HDMapNet. The final objective function combines the loss of diffusion model $\mathcal{L}_{\text{diff}}$ with the auxiliary losses from the baseline model $\mathcal{L}_{\text{baseline}}$, resulting in the following mathematical expression:
@@ III-C2 Inference
As for the sampling in the denoising iterations, we adopt sampler proposed in [29] to accelerate sampling and support the two branch design. Specifically, we can generate a sample $\mathbf{x}_{t-1}$ from a sample $\mathbf{x}_{t}$ via:
@@ IV-A Implementation Details
TABLE IV: Quantitative analysis in different conditions compared with HDMapNet (C+L).
Method Condition Divider Ped. Boundary mIOU Divider Ped. Boundary mAP
HDMapNet* Cloudy $54.7$ $32.7$ $61.6$ $49.7$ $29.8$ $26.9$ 55.1 $37.3$
+DiffMap Cloudy 58.0 35.0 63.3 52.1 35.6 30.3 $52.2$ 39.4
HDMapNet* Rainy $47.7$ $27.8$ $49.6$ $41.7$ $27.6$ $25.1$ 52.3 $35.0$
+DiffMap Rainy 50.4 28.5 49.2 42.7 30.7 27.9 $47.2$ 35.3
HDMapNet* Night $50.5$ 20.6 $60.7$ 43.9 $27.5$ $8.2$ 62.0 $32.6$
+DiffMap Night 53.7 $14.5$ 61.0 $43.1$ 30.4 11.5 $56.8$ 32.9
@@ IV-A1 Dataset
We validate our DiffMap on the nuScenes dataset[30]. As a popular benchmark in autonomous driving, the nuScenes dataset includes multi-view images and point clouds of 1,000 scenes, with 700 scenes allocated for training, 150 for validation, and 150 for testing. The nuScenes dataset also has annotated HD map semantic labels, which are highly suitable for our task as well. Besides training the main model, we rasterize the ego map data into segmentation maps during the pre-training of the autoencoder.
@@ IV-A2 Architecture
We use ResNet-101[31] as the backbone for our camera branch and PointPillars[32] as our LiDAR branch backbone. The Segmentation head in baseline is a FCN network based on ResNet-18. For the autoencoder, we adopt VQVAE, which is pre-trained on the nuScenes segmentation maps dataset, to extract map features and compress map into the basic latent space. UNet is used to construct the diffusion network.